# DOMAIN: milDly cOnservative Model-bAsed offlINe reinforcement learning

### v3

@@ A-B Proof of Theorem 2
$\lambda\cdot\min_{s}\mathbb{E}_{a\sim\pi(a\mid s)}[\eta(s,a)]$ (27)
$\geq$ $\max_{s}\mathbb{E}_{a\sim\pi(a\mid s)}\left[(1-f)\Delta_{1}(s,a)+f\Delta_{2}(s,a)\right]$
Since $\Delta_{1}(s,a)>0,\Delta_{2}(s,a)>0,\lambda>0$, in order for Eq. (27) to satisfy the condition, it is a prerequisite that the term $\min\mathbb{E}_{a\sim\pi(a\mid s)}[\eta(s,a)]>0$ holds, leading to:
$\mathbb{E}_{a\sim\pi(a\mid s)}[\eta(s,a)]$ (28)
$=$ $\mathbb{E}_{a\sim\pi(a\mid s)}\left[\frac{\omega(s,a)-d(s,a)}{(1-f)d^{\pi_{b}}(s,a)+fd_{\widehat{\mathcal{M}}}^{\pi}(s,a)}\right]$
$=$ $\sum_{a}\frac{\omega(s,a)-d(s,a)}{(1-f)d^{\pi_{b}}(s)\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}+fd_{\mathcal{\widehat{\mathcal{M}}}}^{\pi}(s)}$
$\geq$ $\frac{\sum_{a}\omega(s,a)}{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}$
$-\frac{\sum_{a}d(s,a)}{(1-f)d^{\pi_{b}}(s)\min_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{\mathcal{M}}}}^{\pi}(s)}$
@@
$\sum_{a}\omega(s,a)$
$>$ $\frac{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}{(1-f)d^{\pi_{b}}(s)\min_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}\sum_{a}d(s,a)$
$=$ $\xi(s)\sum_{a}d(s,a)$
Substituting Eq. (A-B) into (27), it can further derive:
$\lambda\cdot\min_{s}\frac{(\xi(s)-1)\sum_{a}d(s,a)}{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}$
$\geq$ $(1-f)\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\min_{s,a}\{\sqrt{|\mathcal{D}(s,a)|}\}}$
$+f\left[\max_{s,a}\left\{\left|r_{\mathcal{M}}-r_{\widehat{\mathcal{M}}}\right|\right\}+\frac{2\gamma R_{\max}}{1-\gamma}\max_{s,a}\left\{D\left(T_{\widehat{\mathcal{M}}},T_{\mathcal{M}}\right)\right\}\right]$
$\Rightarrow$ $\lambda\geq\frac{(1-f)\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\min_{s,a}\{\sqrt{|\mathcal{D}|}\}}}{\min_{s}\left\{\frac{(\xi(s)-1)\sum_{a}d(s,a)}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}$
Thus, Theorem 2 is proven.
A-C Proof of Theorem 3
Yu et al. [21] give expression: $\widehat{Q}^{k+1}(s,a)=(\widehat{\mathcal{T}}^{\pi}\widehat{Q}^{k})(s,a)-\lambda[\frac{\rho(s,a)-d(s,a)}{(1-f)d(s,a)+f\rho(s,a)}]$. Assuming DOMAIN and COMBO methods have the same sampling and model errors, we have:
$\kappa_{1}-\kappa_{2}$
$=$ $\mathbb{E}_{s\sim d(s),a\sim\pi(a\mid s)}\left[\widehat{Q}_{1}^{\pi}(s,a)-\widehat{Q}_{2}^{\pi}(s,a)\right]$
@@ A-D Proof of Theorem 4
$|J(\mathcal{M},\pi)-J(\overline{\mathcal{M}},\pi)|$ (29)
$\leq$ $\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\overline{\mathcal{M}}}^{\pi}(s,a)}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$
Proof. The proof follows a similar approach to Kumar et al. [13]. Firstly, we employ the triangle inequality to separate the reward return and the dynamic properties:
$|J(\mathcal{M},\pi)-J(\overline{\mathcal{M}},\pi)|$ (30)
@@
$\leq$ $\frac{1}{1-\gamma}\left|\sum_{s,a}d_{\overline{\mathcal{M}}}^{\pi}(s)\pi(a|s)\big{[}r_{\overline{\mathcal{M}}}(s,a)-r_{\mathcal{M}}(s,a)\big{]}\right|$
$+\frac{1}{1-\gamma}\left|\sum_{s,a}\left(d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right)\pi(a|s)r_{\mathcal{M}}(s,a)\right|$
@@
$\left|\sum_{s,a}d_{\overline{\mathcal{M}}}^{\pi}\pi\left[r_{\overline{\mathcal{M}}}-r_{\mathcal{M}}\right]\right|\leq\mathbb{E}_{(s,a)\sim d_{\overline{\mathcal{M}}}^{\pi}}\left[\frac{C_{r,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$ (31)
@@
$\left|d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right|$ (32)
$\leq$ $\sum_{s,s^{\prime}}\left|\int_{\mathcal{A}}\left(T_{\overline{\mathcal{M}}}-T_{\mathcal{M}}\right)\pi\right|d_{\overline{\mathcal{M}}}^{\pi}(s)$
From Assumption 1, it can deduce that $\left|d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right|\leq\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi}(s,a)}\left[\frac{C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$, resulting in an upper bound for the second term in Eq. (30):
$\left|\sum_{s,a}\left[d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right]\pi(a|s)r_{\mathcal{M}}(s,a)\right|$ (33)
$\leq$ $\mathbb{E}_{(s,a)\sim d_{\mathcal{\overline{\mathcal{M}}}}^{\pi}(s,a)}\left[\frac{R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$
@@
$|J(\mathcal{M},\pi)-J(\overline{\mathcal{M}},\pi)|$ (34)
$\leq$ $\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\overline{\mathcal{M}}}^{\pi}(s,a)}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$
@@
$|J(\mathcal{M},\pi)-J(\widehat{\mathcal{M}},\pi)|$ (35)
$\leq$ $\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\widehat{\mathcal{M}}}^{\pi}(s,a)}\left[\varepsilon_{r}(s,a)+R_{\max}D(s,a)\right]$
Proof. The proof follows a similar process as in Lemma 2. First, it can deduce:
$|J(\mathcal{M},\pi)-J(\widehat{\mathcal{M}},\pi)|$ (36)
$\leq$ $\frac{1}{1-\gamma}\left|\sum_{s,a}d_{\widehat{\mathcal{M}}}^{\pi}(s)\pi(a\mid s)\left[r_{\widehat{\mathcal{M}}}(s,a)-r_{\mathcal{M}}\right]\right|$
$+\frac{1}{1-\gamma}\left|\sum_{s,a}\left(d_{\widehat{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right)\pi(a|s)r_{\mathcal{M}}\right|$
@@
$|J(\mathcal{M},\pi)-J(\widehat{\mathcal{M}},\pi)|$ (37)
$\leq$ $\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\widehat{\mathcal{M}}}^{\pi}(s,a)}\left[\varepsilon_{r}(s,a)+R_{\max}D(s,a)\right]$
Lemma 3 is thus proven. Let $\widehat{Q}^{\pi}(s,a)$ denote the fixed point of Eq. (14). The policy update for the DOMAIN method is expressed as follows:
$\pi^{*}=\max_{\pi}\left[J\left(\mathcal{M}_{f},\pi\right)-\frac{\lambda}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\mathcal{M}_{f}}^{\pi}}[\eta(s,a)]\right]$ (38)
where $J\left(\mathcal{M}_{f},\pi\right)=(1-f)J(\overline{\mathcal{M}},\pi)+fJ(\widehat{\mathcal{M}},\pi)$, $\mathcal{M}_{f}:=(1-f)\overline{\mathcal{M}}+f\widehat{\mathcal{M}}$, and $\eta(s,a)=(\omega(s,a)-d(s,a))/d_{\beta}(s,a)$. Since $\pi^{*}$ represents the optimal value in Eq. (38), and let $\varpi(\pi,f)=\mathbb{E}_{(s,a)\sim d_{\mathcal{M}_{f}}^{\pi}}[\eta(s,a)]$, we have:
$J\left(\mathcal{M}_{f},\pi^{*}\right)-\frac{\lambda\varpi\left(\pi^{*},f\right)}{1-\gamma}\geq J\left(\mathcal{M}_{f},\pi_{b}\right)-\frac{\lambda\varpi\left(\pi_{b},f\right)}{1-\gamma}$ (39)
@@
$\left|J\left(\mathcal{M}_{f},\pi\right)-J(\mathcal{M},\pi)\right|$
$=$ $|(1-f)J(\overline{\mathcal{M}},\pi)+fJ(\widehat{\mathcal{M}},\pi)-J(\mathcal{M},\pi)|$
@@
$\left|J\left(\mathcal{M}_{f},\pi\right)-J(\mathcal{M},\pi)\right|$ (40)
@@
$\leq$ $\frac{1-f}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi}(s,a)}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$
$+\frac{f}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi}(s,a)}\left[\varepsilon_{r}(s,a)+R_{\max}D(s,a)\right]$
Substituting Eq. (40) into (39), it can further derive:
$J\left(\mathcal{M},\pi^{*}\right)-J\left(\mathcal{M},\pi_{b}\right)$ (41)
$\geq$ $\frac{\lambda}{1-\gamma}\left[\varpi\left(\pi^{*},f\right)-\varpi\left(\pi_{b},f\right)\right]$
$-\frac{1-f}{1-\gamma}\left\{\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi^{*}}}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]\right.$
$\left.+\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi_{b}}}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]\right\}$
$-\frac{f}{1-\gamma}\left\{\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi^{*}}}\left[\varepsilon_{r}(s,a)+R_{\max}D(s,a)\right]\right.$
$\left.+\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi_{b}}}\left[\varepsilon_{r}(s,a)+R_{\max}D(s,a)\right]\right\}$
The behavioral policy $\pi_{b}$ is obtained from the offline dataset $\mathcal{D}_{\text{offline }}$, resulting in a larger $\sqrt{|\mathcal{D}(s,a)|}$ under the distribution $d_{\overline{\mathcal{M}}}^{\pi_{b}}$, indicating smaller sampling errors. Additionally, since the environment model is trained using $\mathcal{D}_{\text{offline }}$, the model error is smaller under the distribution $d_{\widehat{\mathcal{M}}}^{\pi_{b}}$. Therefore, Eq. (41) can be further simplified as:
$J\left(\mathcal{M},\pi^{*}\right)-J\left(\mathcal{M},\pi_{b}\right)$ (42)
$\geq$ $\frac{\lambda}{1-\gamma}\left[\varpi\left(\pi^{*},f\right)-\varpi\left(\pi_{b},f\right)\right]$
$-2\frac{1-f}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi^{*}}}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$
$-2\frac{f}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi^{*}}}\Big{[}\varepsilon_{r}(s,a)+R_{\max}D(s,a)\Big{]}$
Theorem 4 is proven.

### v4

@@ VII-B Proof of Theorem 2
$\lambda\cdot\min_{s}\mathbb{E}_{a\sim\pi(a\mid s)}[\eta(s,a)]\geq\max_{s}\mathbb{E}_{a\sim\pi(a\mid s)}\left[(1-f)\Delta_{1}(s,a)+f\Delta_{2}(s,a)\right]$ (S.10)
Since $\Delta_{1}(s,a)>0,\Delta_{2}(s,a)>0$ and $\lambda>0$, for Eq. (S.10) to satisfy the given condition, it is necessary that the term $\min\mathbb{E}_{a\sim\pi(a\mid s)}[\eta(s,a)]$ is strictly positive, which leads to:
$\mathbb{E}_{a\sim\pi(a\mid s)}[\eta(s,a)]=$ $\mathbb{E}_{a\sim\pi(a\mid s)}\left[\frac{\omega(s,a)-d(s,a)}{(1-f)d^{\pi_{b}}(s,a)+fd_{\widehat{\mathcal{M}}}^{\pi}(s,a)}\right]=\sum_{a}\frac{\omega(s,a)-d(s,a)}{(1-f)d^{\pi_{b}}(s)\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}+fd_{\mathcal{\widehat{\mathcal{M}}}}^{\pi}(s)}$ (S.11)
$\geq$ $\frac{\sum_{a}\omega(s,a)}{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}-\frac{\sum_{a}d(s,a)}{(1-f)d^{\pi_{b}}(s)\min_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{\mathcal{M}}}}^{\pi}(s)}$
@@
$\sum_{a}\omega(s,a)>\frac{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}{(1-f)d^{\pi_{b}}(s)\min_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}\sum_{a}d(s,a)=\xi(s)\sum_{a}d(s,a)$ (S.12)
By substituting Eq. (S.12) into (S.10), we can further derive the following:
$\lambda\cdot\min_{s}\frac{(\xi(s)-1)\sum_{a}d(s,a)}{(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)}$ (S.13)
$\geq$ $(1-f)\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\min_{s,a}\{\sqrt{|\mathcal{D}(s,a)|}\}}+f\left[\max_{s,a}\left\{\left|r_{\mathcal{M}}-r_{\widehat{\mathcal{M}}}\right|\right\}+\frac{2\gamma R_{\max}}{1-\gamma}\max_{s,a}\left\{D\left(T_{\widehat{\mathcal{M}}},T_{\mathcal{M}}\right)\right\}\right]$
$\Rightarrow$ $\lambda\geq\frac{(1-f)\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\min_{s,a}\{\sqrt{|\mathcal{D}|}\}}}{\min_{s}\left\{\frac{(\xi(s)-1)\sum_{a}d(s,a)}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}+\frac{f[\max\limits_{s,a}\left\{\left|r_{\mathcal{M}}-r_{\mathcal{M}}\right|\right\}+\frac{2\gamma R_{\max}}{1-\gamma}\max\limits_{s,a}\left\{D_{TV}\left(T_{\mathcal{M}},T_{\widehat{\mathcal{M}}}\right)\right\}]}{\min_{s}\left\{\frac{(\xi(s)-1)\sum_{a}d(s,a)}{\left[(1-f)d^{\pi_{b}}(s)\max_{a}\left\{\frac{\pi_{b}(a\mid s)}{\pi(a\mid s)}\right\}+fd_{\widehat{\mathcal{M}}}^{\pi}(s)\right]}\right\}}$
This completes the proof of Theorem 2.
VII-C Proof of Theorem 3
Following the previous work [28], they give the expression: $\widehat{Q}^{k+1}(s,a)=(\widehat{\mathcal{T}}^{\pi}\widehat{Q}^{k})(s,a)-\lambda[\frac{\rho(s,a)-d(s,a)}{(1-f)d(s,a)+f\rho(s,a)}]$. Assuming DOMAIN and COMBO methods have the same sampling and model errors, we have:
$\kappa_{1}-\kappa_{2}=$ $\mathbb{E}_{s\sim d(s),a\sim\pi(a\mid s)}\left[\widehat{Q}_{1}^{\pi}(s,a)-\widehat{Q}_{2}^{\pi}(s,a)\right]$ (S.14)
@@ VII-D Proof of Theorem 4
$|J(\mathcal{M},\pi)-J(\overline{\mathcal{M}},\pi)|\leq\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\overline{\mathcal{M}}}^{\pi}(s,a)}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$ (S.15)
Proof for Lemma 2. The proof follows a methodology similar to that of Kumar et al. [20]. First, we apply the triangle inequality to decouple the reward return from the dynamic properties:
$|J(\mathcal{M},\pi)-J(\overline{\mathcal{M}},\pi)|$ (S.16)
@@
$\leq$ $\frac{1}{1-\gamma}\left|\sum_{s,a}d_{\overline{\mathcal{M}}}^{\pi}(s)\pi(a|s)\big{[}r_{\overline{\mathcal{M}}}(s,a)-r_{\mathcal{M}}(s,a)\big{]}\right|+\frac{1}{1-\gamma}\left|\sum_{s,a}\left(d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right)\pi(a|s)r_{\mathcal{M}}(s,a)\right|$
@@
$\left|\sum_{s,a}d_{\overline{\mathcal{M}}}^{\pi}\pi\left[r_{\overline{\mathcal{M}}}-r_{\mathcal{M}}\right]\right|\leq\mathbb{E}_{(s,a)\sim d_{\overline{\mathcal{M}}}^{\pi}}\left[\frac{C_{r,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$ (S.17)
@@
$\left|d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right|\leq$ $\sum_{s,s^{\prime}}\left|\int_{\mathcal{A}}\left(T_{\overline{\mathcal{M}}}-T_{\mathcal{M}}\right)\pi\right|d_{\overline{\mathcal{M}}}^{\pi}(s)$ (S.18)
From Assumption 1, it can be deduced that
$\left|d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right|\leq\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi}(s,a)}\left[\frac{C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$ (S.19)
This result establishes an upper bound for the second term in Eq. (S.16):
$\left|\sum_{s,a}\left[d_{\overline{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right]\pi(a|s)r_{\mathcal{M}}(s,a)\right|\leq\mathbb{E}_{(s,a)\sim d_{\mathcal{\overline{\mathcal{M}}}}^{\pi}(s,a)}\left[\frac{R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$ (S.20)
@@
$|J(\mathcal{M},\pi)-J(\overline{\mathcal{M}},\pi)|\leq\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\overline{\mathcal{M}}}^{\pi}(s,a)}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]$ (S.21)
@@
$|J(\mathcal{M},\pi)-J(\widehat{\mathcal{M}},\pi)|\leq\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\widehat{\mathcal{M}}}^{\pi}(s,a)}\left[\varepsilon_{r}(s,a)+R_{\max}D(s,a)\right]$ (S.22)
Proof for Lemma 3. The proof follows a similar process as in Lemma 2. First, it can deduce:
$|J(\mathcal{M},\pi)-J(\widehat{\mathcal{M}},\pi)|$ (S.23)
$\leq$ $\frac{1}{1-\gamma}\left|\sum_{s,a}d_{\widehat{\mathcal{M}}}^{\pi}(s)\pi(a\mid s)\left[r_{\widehat{\mathcal{M}}}(s,a)-r_{\mathcal{M}}\right]\right|+\frac{1}{1-\gamma}\left|\sum_{s,a}\left(d_{\widehat{\mathcal{M}}}^{\pi}(s)-d_{\mathcal{M}}^{\pi}(s)\right)\pi(a|s)r_{\mathcal{M}}\right|$
@@
$|J(\mathcal{M},\pi)-J(\widehat{\mathcal{M}},\pi)|\leq\frac{1}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\widehat{\mathcal{M}}}^{\pi}(s,a)}\left[\varepsilon_{r}(s,a)+R_{\max}D(s,a)\right]$ (S.24)
Lemma 3 is thus proven.
Proof for Theorem 4. Let $\widehat{Q}^{\pi}(s,a)$ denote the fixed point of Eq. (14). The policy update for the DOMAIN method is expressed as follows:
$\pi^{*}=\max_{\pi}\left[J\left(\mathcal{M}_{f},\pi\right)-\frac{\lambda}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\mathcal{M}_{f}}^{\pi}}[\eta(s,a)]\right]$ (S.25)
where $J\left(\mathcal{M}_{f},\pi\right)=(1-f)J(\overline{\mathcal{M}},\pi)+fJ(\widehat{\mathcal{M}},\pi)$, $\mathcal{M}_{f}:=(1-f)\overline{\mathcal{M}}+f\widehat{\mathcal{M}}$, and $\eta(s,a)=(\omega(s,a)-d(s,a))/d_{\beta}(s,a)$. Since $\pi^{*}$ represents the optimal value in Eq. (S.25), and let $\varpi(\pi,f)=\mathbb{E}_{(s,a)\sim d_{\mathcal{M}_{f}}^{\pi}}[\eta(s,a)]$, we have:
$J\left(\mathcal{M}_{f},\pi^{*}\right)-\frac{\lambda\varpi\left(\pi^{*},f\right)}{1-\gamma}\geq J\left(\mathcal{M}_{f},\pi_{b}\right)-\frac{\lambda\varpi\left(\pi_{b},f\right)}{1-\gamma}$ (S.26)
@@
$\left|J\left(\mathcal{M}_{f},\pi\right)-J(\mathcal{M},\pi)\right|=$ $|(1-f)J(\overline{\mathcal{M}},\pi)+fJ(\widehat{\mathcal{M}},\pi)-J(\mathcal{M},\pi)|$ (S.27)
@@
$\left|J\left(\mathcal{M}_{f},\pi\right)-J(\mathcal{M},\pi)\right|$ (S.28)
@@
$\leq$ $\frac{1-f}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi}(s,a)}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]+\frac{f}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi}(s,a)}\left[\varepsilon_{r}(s,a)+R_{\max}D(s,a)\right]$
Substituting Eq. (S.28) into (S.26), it can further derive:
$J\left(\mathcal{M},\pi^{*}\right)-J\left(\mathcal{M},\pi_{b}\right)$ (S.29)
$\geq$ $\frac{\lambda}{1-\gamma}\left[\varpi\left(\pi^{*},f\right)-\varpi\left(\pi_{b},f\right)\right]-\frac{1-f}{1-\gamma}\left\{\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi^{*}}}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]\right.\left.+\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi_{b}}}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]\right\}$
$-\frac{f}{1-\gamma}\left\{\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi^{*}}}\left[\varepsilon_{r}(s,a)+R_{\max}D(s,a)\right]\right.\left.+\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi_{b}}}\left[\varepsilon_{r}(s,a)+R_{\max}D(s,a)\right]\right\}$
The behavioral policy $\pi_{b}$ is derived from the offline dataset $\mathcal{D}_{\text{offline }}$, resulting in a larger $\sqrt{|\mathcal{D}(s,a)|}$ under the distribution $d_{\overline{\mathcal{M}}}^{\pi_{b}}$, which indicates smaller sampling errors. Furthermore, since the environment model is trained using $\mathcal{D}_{\text{offline }}$, the model error is reduced under the distribution $d_{\widehat{\mathcal{M}}}^{\pi_{b}}$. Therefore, Eq. (S.29) can be further simplified as:
$J\left(\mathcal{M},\pi^{*}\right)-J\left(\mathcal{M},\pi_{b}\right)$ (S.30)
$\geq$ $\frac{\lambda}{1-\gamma}\left[\varpi\left(\pi^{*},f\right)-\varpi\left(\pi_{b},f\right)\right]-2\frac{1-f}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi^{*}}}\left[\frac{C_{r,\delta}+R_{\max}C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}\right]-2\frac{f}{1-\gamma}\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi^{*}}}\Big{[}\varepsilon_{r}(s,a)+R_{\max}D(s,a)\Big{]}$
This completes the proof of Theorem 4.